OpenStudy (anonymous):

What is the solution to the equation one third (9x-6)=2x+3? X=

OpenStudy (anonymous):

Something like this? \[\frac{ 9x-6 }{ 3 }=2x+3\]

OpenStudy (anonymous):

\[\frac{ 1}{ 3} (9x-6)=2x+3\]

OpenStudy (anonymous):

That is what I wrote...anyway we need to simplify the left side of the equation. How would you do that?

OpenStudy (anonymous):

the Parentheses' first and then the rest

OpenStudy (anonymous):

How do we get rid of the 3? \[1/3 (9x-6) = \frac{ 9x-6 }{ 3 }\]

OpenStudy (anonymous):

\[\frac{ 9x -6 }{ 3 }= What\]

OpenStudy (anonymous):

No you do not understand. 1/3 times 9x -6 is the same as 9x-6 /3.\[\frac{ 1 }{ 3 }x \frac{ 9x-6 }{ 1 } = \frac{ 9x -6 }{ 3 }\]

OpenStudy (anonymous):

oh okay sorry

OpenStudy (anonymous):

No problemo. So how do you simplify 9x-6 /3?

OpenStudy (anonymous):

Divide each term above the line by 3. So 9x divided by 3 equals what? And -6 divided by 3 equals what?

OpenStudy (anonymous):

3x and -2

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

First we simplify: \(\dfrac{1}{3}(9x - 6)\) Can you multiply these? \(\dfrac{1}{3} \times 9x\) \(\dfrac{1}{3} \times -6\)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

3 and 2

OpenStudy (igreen):

Actually that would be 3x and -2.

OpenStudy (igreen):

So we have: \(3x - 2 = 2x + 3\) Add 2 to both sides: \(3x = 2x + 5\) Subtract 2 to both sides, what's 3x - 2x? @Alpha_Female_Wolf02